OpenStudy (anonymous):

the slope of a line passing though (-5,10) and (-9,3) is?

jimthompson5910 (jim_thompson5910):

Two generic points that define a straight line are (x1, y1) and (x2, y2) In this case, (x1, y1) = (-5, 10), which means x1=-5 and y1=10 and (x2, y2) = (-9, 3), which means x2=-9 and y2=3 So x1=-5, y1=10, x2=-9, and y2=3. ------------------- Now use the slope formula (and plug those values in found earlier) m = (y2-y1)/(x2-x1) m = (3-10)/(-9-(-5)) m = (3-10)/(-9+5) m = (-7)/(-4) m = 7/4 So the slope of the line that passes through the two points is 7/4